if a rectangle has an area of 875 sq ft what is its length, perimeter, width
Answer by Earlsdon(6103)

It's impossible to say with the limited information provided.
A = L*W
875 = L*W
--------------
875 = 1*875
875 = 5*175
875 = 10*87.5
875 = 25*35
875 = 7*125
Need I go on?